OpenStudy (anonymous):

find the x-intercepts and y-intercepts f(x)= -x^2+2x+8

OpenStudy (anonymous):

An x intercept is when y = 0, and a y intercept is when x =0. Plug 0 in for x to find the y-intercept. To find the x intercepts, plug your a (-1) b (2) and c (8) values into the quadratric formula

OpenStudy (anonymous):

f(x)= -x^2+2x+8 , for x-intercept let -x^2+2x+8=0 ==> -( x^2-2x-8)=0 ===> (x-4)(x+2)=0 ===> x=4 or x=-2

OpenStudy (anonymous):

for the y-intercept let x=0 ===> y=0+0+8 =8